# Demystifying Invariant Effectiveness for Securing Smart Contracts

### v1

@@ Abstract.
Smart contract transactions associated with security attacks often exhibit distinct behavioral patterns compared with historical benign transactions before the attacking events. While many runtime monitoring and guarding mechanisms have been proposed to validate invariants and stop anomalous transactions on the fly, the empirical effectiveness of the invariants used remains largely unexplored. In this paper, we studied $23$ prevalent invariants of $8$ categories, which are either deployed in high-profile protocols or endorsed by leading auditing firms and security experts. Using these well-established invariants as templates, we developed a tool Trace2Inv which dynamically generates new invariants customized for a given contract based on its historical transaction data. We evaluated Trace2Inv on $42$ smart contracts that fell victim to $27$ distinct exploits on the Ethereum blockchain. Our findings reveal that the most effective invariant guard alone can successfully block $18$ of the $27$ identified exploits with minimal gas overhead. Our analysis also shows that most of the invariants remain effective even when the experienced attackers attempt to bypass them. Additionally, we studied the possibility of combining multiple invariant guards, resulting in blocking up to $23$ of the $27$ benchmark exploits and achieving false positive rates as low as $0.32\%$. Trace2Inv outperforms current state-of-the-art works on smart contract invariant mining and transaction attack detection in terms of both practicality and accuracy. Though Trace2Inv is not primarily designed for transaction attack detection, it surprisingly found two previously unreported exploit transactions, earlier than any reported exploit transactions against the same victim contracts.
copyright: nonecopyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXbooktitle: Companion Proceedings of the 32nd ACM Symposium on the Foundations of Software Engineering (FSE ’24), July 15–19, 2024, Porto de Galinhas, Brazilprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Blockchain technology has paved the way for decentralized, resilient, and programmable ledgers on a global scale. One of its most impactful applications is smart contracts, which can be deployed onto a blockchain. These smart contracts allow developers to encode intricate transaction rules that govern the ledger. This innovation has made both blockchains and smart contracts essential infrastructure for decentralized financial services, commonly known as DeFi. As of Sept 25, 2023, the Total Value Locked (TVL) in $2,933$ DeFi protocols has reached an impressive $48.58$ billion (def, 2023a).
@@
However, the landscape is not without its challenges. Security attacks pose a significant threat to the security of smart contracts. Attackers can exploit various vulnerabilities by sending malicious transactions, potentially leading to the theft of millions of dollars from users. As of Sept 25, 2023, the financial losses attributed to security attacks on DeFi protocols exceeded $5.53$ billion USD (def, 2023b).
@@
One key observation is that transactions initiated by attackers often display abnormal behaviors when compared to standard transactions from regular DeFi contract users. These malicious transactions may exploit control flows in corner cases, use abnormally large values to trigger overflows, or manipulate a large volume of digital assets to distort the market in DeFi contracts. In fact, industry experts have been actively monitoring abnormal digital asset movements on-chain to report malicious activities. For example, Forta Network (Network, 2023) deploys monitoring bots to detect on-chain security-related events in real-time. Driven by this observation, smart contract developers have proposed deploying runtime checks to detect transactions leading to abnormal behaviors to neutralize malicious attacks. These checks involve enforcing various runtime invariants, such as restricting the maximum number of digital asset deposits or withdrawals in a contract to prevent market manipulation. Another example is to limit the interaction of other contracts to prevent attackers from crafting sophisticated attack strategies. However, these mechanisms are often manually designed and tailored for specific contracts. This raises questions about their effectiveness across different types of contracts and whether they maintain an acceptable false positive rate without hindering normal user activities.
@@
Smart Contract Invariant Study: This paper presents the first comprehensive, quantitative analysis focused on the utilization of dynamically inferred invariants to enhance smart contract security. We examine $23$ invariant templates, which are advocated by leading auditing firms, academic research, and DeFi protocol developers. Our findings indicate that dynamically inferred dynamic invariants serve as effective mechanisms for thwarting security breaches. When appropriately configured, these invariants can neutralize over 74.1% of malicious attacks while maintaining a false positive rate of less than 0.32%.
@@ 2. Background
Blockchain is a distributed, immutable ledger technology that records transactions across multiple nodes in a network. It employs cryptographic techniques to ensure data integrity and consensus algorithms to maintain final consistency across all participating nodes. Smart contracts are self-executing contracts with the terms of the agreement directly written into code. Deployed on blockchains, they are immutable and transparent, enabling trustless transactions without the need for intermediaries. Invariant guards (also called circuit breakers) are runtime checks around contract invariant conditions that shall always hold during contract exeuction, aiming to secure smart contracts on the fly.
Ethereum Virtual Machine (EVM) is the runtime environment for smart contract execution on Ethereum. It is a Turing-complete virtual machine that interprets and executes the bytecode compiled from contracts programmed in high-level language like Solidity. Gas is a unit of transaction fee on Ethereum, used to quantify the computational efforts for the execution of EVM operations. It is paid in Ether, the native cryptocurrency of Ethereum. Externally Owned Account (EOA) and contract account are two types of accounts on Ethereum. EOAs are owned by normal users only who have the right to send transactions to blockchains, while contract accounts are controlled by the code deployed at a certain blockchain address and its code will be executed when the contract function is invoked. ERC20 is a standard interface for fungible tokens on Ethereum. Almost all valuable tokens on Ethereum are ERC20 tokens.
Common smart contract vulnerabilities include integer overflow/underflow (Blockchain-Projects, 2020), reentrancy (Siegel, 2016), dangerous delegatecall (Santiago, 2017), etc. DeFi vulnerabilities are smart contract vulnerabilities that are specific to DeFi applications. They are more subtle to detect and attacks usually involve more sophisticated steps (Zhou et al., 2023). DeFi protocols are major targets for smart contract attacks, which experienced \$5.53 billion loss out of the overall \$6.94 billion in recent blockchain incidents (def, 2023b).
@@ 3. Motivating Example
Specifically, the attacker in the exploit transaction first borrows a large amount of digital assets and uses the borrowed asset to buy USDC in Curve to inflate the USDC price in Curve. Then it deposits $49.98$M USDC into Harvest vault contract, which increases its USDC balance from $72.83$M to $122.51$M. Due to the manipulated oracle price of USDC, Harvest vault contract erroneously mints the attacker an inflated $51.46$M fUSDC, which increases the total fUSDC supply from $127.58$M to $179.04$M. The attacker then restores the USDC by selling the USDC in Curve, and redeems all its fUSDC tokens for $50.30$M USDC, yielding a $32$k surplus compared to the initial deposit. This redemption decreases the Harvest vault’s USDC balance from $122.81$M to $72.51$M and restores the total fUSDC supply to its original value of $127.58$M. Remarkably, this identical attack vector is executed three times within one exploit transaction, consuming an unusually high gas count of $9,895,111$, narrowly within the gas limit of $12,065,986$ at the time.
@@
RQ5: In terms of enhancing smart contract security, how does this work compare to other state-of-the-art works in smart contract invariant generation or transaction anomaly detection?
@@ 4. Invariants
Methodology. In our effort to build a comprehensive list of smart contract transaction guards, we carried out an analysis of existing research papers on smart contract security (Liu et al., 2022b; Ghaleb et al., 2023; Rodler et al., 2018; Choi et al., 2021; Breidenbach et al., 2021). Additionally, we conducted a qualitative study on both audit reports and source code of the $63$ audited projects from ConsenSys, a leading smart contract auditing firm, from May 2020 to March 2023 (con, 2023). One author extracted $2181$ enforced invariants from the audit reports and smart contracts under audit by searching for keywords such as “require” and “assert”, after eliminating any duplicates. The author then manually reviewed these invariants to extract templates for pattern matching against remaining uncategorized invariants. This iterative process continued until no new templates could be extracted, and all remaining invariants were also deemed uncategorized for specific reasons. This task took three weeks. Following this, another author reviewed and validated both the categorized and uncategorized invariants for accuracy. In cases of disagreement, a third author was consulted to resolve the issue. This review process lasted two weeks. All three authors have over two years of smart contract security research experience.
Table 1. Qualitative Study Statistics Overview (The table’s left section presents key statistics from the qualitative study. The middle section presents categorized instances across invariant categories. The right section presents instances for various reasons why these invariants remain uncategorized.)
@@
The above process resulted in $826$ invariants under $8$ categories, which comprise $37.87\%$ of the overall $2181$ invariants, as shown in Table 1. Access Control and DataFlow are the top and second common categories, respectively. The remaining invariants ($62.13\%$) can not be categorized for various reasons. The most common reason is that invariants are specific to a particular protocol. For example, the invariant “require(validUniswapPath(bAsset))” checks whether “bAsset” is a valid Uniswap path. However, this invariant can only apply to protocols involving Uniswap, thus limiting its applicability. Other uncategorized invariants are used for checking array lengths, byte operations, and arithmetic safety, which target specific data structures or operations. Such invariants of low-level operations are hard to apply as security guards because they are unable to capture the high-level user intentions.
@@
Access Control EOA onlyEOA $\mathsf{msg.sender}=\mathsf{tx.origin}$ (Liu et al., 2022b; Ghaleb et al., 2023)
@@
OD OracleDeviation $|(\mathsf{r_{newPr}}-\mathsf{x_{oldPr}})/\mathsf{x_{oldPr}}|\leq\mathsf{pr_{Dev}?}$ Integer (dfo, 2023b; Breidenbach et al., 2021)
Special Storage TSU TotalSupplyUpperBound $\mathsf{x_{totSup}}\leq\mathsf{totSup?}$ Integer (Aav, 2023; dfo, 2023a)
TBU TotalBorrowUpperBound $\mathsf{x_{totBor}}\leq\mathsf{totBor?}$ Integer (Aav, 2023; dfo, 2023a)
MoneyFlow TIU TokenInUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
@@
TOU TokenOutUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023; dfo, 2023a)
@@
4.1. Access Control
@@
onlyEOA (EOA): This template verifies that the transaction’s origin matches the sender’s address, thereby confirming it was initiated from an externaly owned user address (i.e., EOA address) rather than a contract address. The intuition of this invariant template is that many attack strategies involves multiple sophisticated interactions and therefore attackers often have to write their own contracts. This template can neutralize such attack strategies.
@@ 4.5. Oracle Slippage
The Oracle Slippage category mitigates risks tied to price oracles in decentralized finance (DeFi) applications. The intuition of templates in this category is to detect potential price manipulation by malicious attacks. This class includes two invariant templates: OracleRange (OR) and OracleDeviation (OD). The OR template enforces a bounded range for the oracle prices. It utilizes two parameters, $\mathsf{pr_{LB}?}$ and $\mathsf{pr_{UB}?}$. The OD template enforces a specific percentage deviation limit between the current and last price provided by the oracle. The parameter $\mathsf{pr_{Dev}?}$ is employed to define a permissible deviation rate. These invariants are usually inserted right after the oracle is called.
@@ 4.8. DataFlow
MappingUpperBound (MU): This invariant focuses on values stored in the contract’s mapping data structure, often representing a user’s property(e.g., a user’s share in a liquidity pool). To constrain such user-specific values, we introduce a parameter $\mathsf{v?}$ to set an upper limit on these mappings.
@@
The above invariants are inserted at the locations where data flow variables are first read. This is often in functions that initiate token transfers. These invariants helps the contract ensure that every value used for the calculation of token transfers is within the normal range.
@@ 5. Trace2Inv
We present Trace2Inv, our proposed framework to infer concrete invariants as introduced in Section 4 for a given contract by analyzing its historical transaction traces. As shown in Figure 2, Trace2Inv is structured into three modules: trace parser, invariant-related data extraction, and invariant generation. The second module consists of three sub-modules: invocation tree analysis, type inference, and dynamic taint analysis.

### v2

@@ Abstract.
Smart contract transactions associated with security attacks often exhibit distinct behavioral patterns compared with historical benign transactions before the attacking events. While many runtime monitoring and guarding mechanisms have been proposed to validate invariants and stop anomalous transactions on the fly, the empirical effectiveness of the invariants used remains largely unexplored. In this paper, we studied $23$ prevalent invariants of $8$ categories, which are either deployed in high-profile protocols or endorsed by leading auditing firms and security experts. Using these well-established invariants as templates, we developed a tool Trace2Inv which dynamically generates new invariants customized for a given contract based on its historical transaction data. We evaluated Trace2Inv on $42$ smart contracts that fell victim to $27$ distinct exploits on the Ethereum blockchain. Our findings reveal that the most effective invariant guard alone can successfully block $18$ of the $27$ identified exploits with minimal gas overhead. Our analysis also shows that most of the invariants remain effective even when the experienced attackers attempt to bypass them. Additionally, we studied the possibility of combining multiple invariant guards, resulting in blocking up to $23$ of the $27$ benchmark exploits and achieving false positive rates as low as $0.28\%$. Trace2Inv significantly outperforms state-of-the-art works on smart contract invariant mining and transaction attack detection in accuracy. Trace2Inv also surprisingly found two previously unreported exploit transactions.
runtime validation, invariant generation, dynamic analysis, flash loan
copyright: rightsretaineddoi: 10.1145/3660786journalyear: 2024submissionid: fse24main-p241-pjournal: PACMSEjournalvolume: 1journalnumber: FSEarticle: 79publicationmonth: 7ccs: Security and privacy Software security engineeringccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Blockchain technology has paved the way for decentralized, resilient, and programmable ledgers on a global scale. One of its most impactful applications is smart contracts. These smart contracts allow developers to encode intricate transaction rules that govern the ledger. This innovation has made both blockchains and smart contracts essential infrastructure for decentralized financial services, commonly known as DeFi. As of Sept 25th, 2023, the Total digital asset Value Locked (TVL) in $2,933$ DeFi protocols has reached an impressive $48.58$ billion (def, 2023a).
@@
However, the landscape is not without its challenges. Security attacks pose a significant threat to the security of smart contracts. Attackers can exploit various vulnerabilities by sending malicious transactions, potentially leading to the theft of millions of dollars. As of Sept 25th, 2023, the financial losses attributed to security attacks on DeFi protocols exceeded $5.53$ billion USD (def, 2023b).
@@
One key observation is that transactions initiated by attackers often display abnormal behaviors when compared to standard transactions from regular DeFi contract users. These malicious transactions may exploit control flows in corner cases, use abnormally large values to trigger overflows, or manipulate a large volume of digital assets to distort the market in DeFi contracts. In fact, industry experts have been actively monitoring abnormal digital asset movements on-chain to report malicious activities. For example, Forta Network (Network, 2023) deploys monitoring bots to detect on-chain security-related events in real-time. Driven by this observation, smart contract developers have proposed deploying runtime checks to detect transactions leading to abnormal behaviors to neutralize malicious attacks. These checks involve enforcing various runtime invariants, such as restricting the maximum number of digital asset deposits or withdrawals in a contract to prevent market manipulation. Another example is to limit the interaction between contracts to prevent attackers from crafting sophisticated attack strategies. However, these mechanisms are often manually designed and tailored for specific DeFi protocols. This raises questions about their effectiveness across different types of contracts and whether they maintain an acceptable false positive rate without hindering normal user activities.
@@
Smart Contract Invariant Study: This paper presents the first comprehensive, quantitative analysis focused on the utilization of dynamically inferred invariants to enhance smart contract security. We examine $23$ invariant templates, which are advocated by leading auditing firms, academic research, and DeFi protocol developers. Our findings indicate that dynamically inferred dynamic invariants serve as effective mechanisms for thwarting security breaches. This paper then proposes new strategies to combine multiple invariants effectively. Our combined invariants can neutralize over 74.1% of malicious attacks while maintaining a false positive rate of less than 0.28%.
@@ 2. Background
Blockchain is a distributed and immutable ledger technology that records transactions across multiple nodes in a network. It employs cryptographic techniques to ensure data integrity and consensus algorithms to maintain final consistency across all participating nodes. Smart contracts are self-executing programs with the terms of the agreement directly written into code. Deployed on blockchains, they are immutable and transparent, enabling trustless transactions without the need for intermediaries. Invariant guards (also called circuit breakers) are runtime checks around contract invariant conditions that shall always hold during contract execution, aiming to secure smart contracts on the fly.
Ethereum Virtual Machine (EVM) is the runtime environment for smart contract execution on Ethereum. It is a Turing-complete virtual machine that interprets and executes the bytecode compiled from contracts programmed in a high-level language like Solidity. Gas is a unit of transaction fee on Ethereum, used to quantify the computational efforts for the execution of EVM operations. It is paid in Ether, the native cryptocurrency of Ethereum. Externally Owned Account (EOA) and contract account are two types of accounts on Ethereum. EOAs are owned by normal users only who have the right to send transactions to blockchains, while contract accounts are controlled by the code deployed at a certain blockchain address and its code will be executed when the contract function is invoked. ERC20 is a standard interface for fungible tokens on Ethereum. Almost all valuable tokens on Ethereum are ERC20 tokens.
Common smart contract vulnerabilities include integer overflow/underflow (Blockchain-Projects, 2020), reentrancy (Siegel, 2016), dangerous delegatecall (Santiago, 2017), etc. DeFi vulnerabilities are smart contract vulnerabilities that are specific to DeFi applications. They are more subtle to detect and attacks usually involve more sophisticated steps (Zhou et al., 2023). DeFi protocols are major targets for smart contract attacks, which have experienced \$5.53 billion loss out of the overall \$6.94 billion caused by recent blockchain incidents (def, 2023b).
@@ 3. Motivating Example
Specifically, the attack transaction first borrows a large amount of digital assets and uses the borrowed asset to buy USDC in Curve to inflate its price. Then it deposits $49.98$M USDC into Harvest vault contract, which increases its USDC balance from $72.83$M to $122.51$M. Due to the manipulated oracle price of USDC, Harvest vault contract erroneously mints the attacker an inflated $51.46$M fUSDC, which increases the total fUSDC supply from $127.58$M to $179.04$M. The attacker then restores the USDC by selling the USDC in Curve, and redeems all its fUSDC tokens for $50.30$M USDC, yielding a $32$k surplus compared to the initial deposit. This redemption decreases the Harvest vault’s USDC balance from $122.81$M to $72.51$M and restores the total fUSDC supply to its original value of $127.58$M. Remarkably, this identical attack vector is executed three times within one exploit transaction, consuming an unusually high gas count of $9,895,111$, narrowly within the gas limit of $12,065,986$ at the time. More details of this example can be found at the website (Chen et al., 2024e).
@@
We observed several other outlier data points besides the last data point. For access frequency, two blocks contained two withdraw invocations within the same block, which, upon investigation, were found to originate from separate addresses. We believe it was a coincidence that two users withdrew their funds in the same block. For token flow, a very large USDC deposit occurred immediately after the contract’s deployment, which, upon manual analysis, was identified as an initial liquidity provided by the Harvest Finance team. Similarly, the abrupt jump in the total supply of fUSDC tokens in the early stage was also due to this initial liquidity provision. Since users receive fUSDC tokens when they deposit USDC into the vault and these tokens are burned upon withdrawal of USDC, a large initial deposit would lead to a corresponding increase in the total supply of fUSDC tokens. For gas consumption, a transaction at block 11094671 consumed 5,062,934 gas, which, after manual analysis, was found to be sent by the Harvest Finance team for migrating old vaults to new ones. This migration process involved numerous operations in a single transaction to ensure atomicity, leading to high gas consumption. Overall, while there were anomalies during the early stage of the contract’s deployment, the different transaction metrics stabilized after some blocks until significantly changed by the exploit transaction. If contract maintainers can identify the ”stable” stage and use these transactions to infer invariants, the inferred invariants can be much more accurate. However, in this paper, we assume such information is not available.
@@
RQ5: In terms of enhancing smart contract security, how does this work compare to other state-of-the-art works in contract invariant generation or transaction anomaly detection?
@@ 4. Invariants
Methodology. In our effort to build a comprehensive list of smart contract transaction guards, we carried out an analysis of existing research papers on smart contract security (Liu et al., 2022b; Ghaleb et al., 2023; Rodler et al., 2018; Choi et al., 2021; Breidenbach et al., 2021).
Additionally, we conducted a qualitative study on both audit reports and source code of the $63$ audited projects from ConsenSys, a leading smart contract auditing firm, from May 2020 to March 2023 (con, 2023). One author extracted $2,181$ enforced invariants from the audit reports and smart contracts under audit by searching for keywords such as “require” and “assert”, after eliminating any duplicates. The author then manually reviewed these invariants to extract templates for pattern matching against remaining uncategorized invariants. This iterative process continued until no new templates could be extracted, and all remaining invariants were also deemed uncategorized for specific reasons. This task took three weeks. Following this, another author reviewed and validated both the categorized and uncategorized invariants for accuracy. In cases of disagreement, a third author was consulted to resolve the issue. This review process lasted two weeks. All three authors have over two years of smart contract security research experience.
Table 1. Qualitative Study Statistics Overview+ (The table’s left section presents key statistics from the qualitative study. The middle section presents categorized instances across invariant categories. The right section presents instances for various reasons why these invariants remain uncategorized.)
@@
The above process resulted in $826$ invariants under $8$ categories, which represent $37.87\%$ of the $2181$ invariants, as shown in Table 1. Access Control and Data Flow are the two common categories. The remaining invariants ($62.13\%$) were not categorized for various reasons. The most common reason is that invariants are specific to a particular protocol. For example, the invariant “require(validUniswapPath(bAsset))” checks whether “bAsset” is a valid Uniswap path. However, this invariant can only apply to protocols involving Uniswap, thus limiting its applicability. Other uncategorized invariants are used for checking array lengths, byte operations, and arithmetic safety, which target specific data structures or operations. Such invariants of low-level operations are hard to apply as security guards because they are unable to capture the high-level user intentions.
@@
Access Control EOA onlyEOA $\mathsf{msg.sender}=\mathsf{tx.origin}$ (Liu et al., 2022b) (Ghaleb et al., 2023)
@@
OD OracleDeviation $|(\mathsf{r_{newPr}}-\mathsf{x_{oldPr}})/\mathsf{x_{oldPr}}|\leq\mathsf{pr_{Dev}?}$ Integer
(dfo, 2023b)
(Breidenbach et al., 2021)
Special Storage TSU TotalSupplyUpperBound $\mathsf{x_{totSup}}\leq\mathsf{totSup?}$ Integer (Aav, 2023) (dfo, 2023a)
TBU TotalBorrowUpperBound $\mathsf{x_{totBor}}\leq\mathsf{totBor?}$ Integer (Aav, 2023) (dfo, 2023a)
Money Flow TIU TokenInUpperBound $\mathsf{r_{tokenIn}}\leq\mathsf{v?}$ Integer (bal, 2023) (dfo, 2023a)
@@
TOU TokenOutUpperBound $\mathsf{r_{tokenOut}}/\mathsf{b_{token,adr}}\leq\mathsf{v?}$ Integer (bal, 2023) (dfo, 2023a)
@@
4.1. Access Control (also called Permission Control)
@@
onlyEOA (EOA) [^1] It is important to note that the forthcoming Spectra upgrade of Ethereum, anticipated for late 2024 or early 2025, will implement EIP-3074 (EIP, 2024b). This implementation is expected to make EOA completely bypassable. This template verifies that the transaction’s origin matches the sender’s address, thereby confirming it was initiated from an externaly owned user address (i.e., EOA address) rather than a contract address. The intuition of this invariant template is that many attack strategies involves multiple sophisticated interactions and therefore attackers often have to write their own contracts. This template can neutralize such attack strategies.
@@ 4.5. Oracle Slippage
The Oracle Slippage category mitigates risks tied to price oracles in DeFi applications. The intuition of templates in this category is to detect potential price manipulation by malicious attacks. This class includes two invariant templates: OracleRange (OR) and OracleDeviation (OD). The OR template enforces a bounded range for the oracle prices. It utilizes two parameters, $\mathsf{pr_{LB}?}$ and $\mathsf{pr_{UB}?}$. The OD template enforces a specific percentage deviation limit between the current and last price provided by the oracle. The parameter $\mathsf{pr_{Dev}?}$ is employed to define a permissible deviation rate. These invariants are usually inserted right after the oracle is called.
@@ 4.8. Data Flow
MappingUpperBound (MU): This invariant focuses on values stored in the contract’s mapping data structure, often representing a user’s property(e.g., a user’s shares/balances). To constrain such user-specific values, we introduce a parameter $\mathsf{v?}$ to set an upper limit on these mappings.
@@
The above invariants are inserted at the locations where data flow variables are first read. This is often in functions that initiate token transfers. These invariants help the contract ensure that every value used for the calculation of token transfers is within the normal range.
@@ 5. Trace2Inv
We present Trace2Inv, shown in Figure 2, a framework to infer concrete invariants as introduced in Section 4 for a contract by analyzing its transaction traces. Trace2Inv consists of three modules: trace parser, invariant-related data extraction, and invariant generation. The second module consists of three submodules: invocation tree analysis, type inference, and dynamic taint analysis.